# Transformation of orientation and rotation angles of synchronous satellites: Application to the Galilean moons

### v1

@@ 6.2 Series for obliquity, node longitude, and offset
The temporal evolution of the inertial ($\theta$, Eq. 11a) and orbital ($\varepsilon$, Eq. 13) obliquities, node longitude ($\psi$, Eq. 11b), and offset ($\zeta$, Eq. 16) obtained from the series of Table 3 and Table 4 are shown in Figs. 3, 4, and 5. We see that the deviation from the Cassini plane is small for Europa (usually less than $0.006^{\circ}$). For comparison with the spin Euler angles, the inclination ($i$, Eq. 12a) and orbital node longitude ($\Omega$, Eq. 12b) are also shown in Fig. 3 and 4, respectively. The typical precision (or truncation error) of the time-series $\varepsilon(t)$ and $\theta(t)$ over the 1900-2100 interval, reconstructed from Table 4, is by construction similar to that of the time-series $s_{x/y}(t)$, that is approximately $0.0003^{\circ}$ (see Table 5). This corresponds to a relative error below $1\%$, except in the case of Io’s orbital obliquity which is particularly small. In contrast, the longitude of the ascending node is a circulating angle, and the error associated with its reconstruction using Table 4 and Eq. (11b) is significantly larger but remains below $0.6^{\circ}$ and therefore below the percent level. The absolute precision on the reconstructed $\zeta(t)$ is similar to that of $\varepsilon(t)$, but its relative precision may be larger than that of $\varepsilon(t)$ when the offset is small (e.g. Europa, $5\%$). As for $s_{x/y}(t)$, neglecting terms related to trends would lead after $100$ years to typical errors larger than the truncation error in $\theta$, $\psi$, and $\zeta$ ($\varepsilon$ is not affected by the trends).
When we inject the series ($\Delta n_{x},\Delta n_{y}$) and ($\Delta s_{x},\Delta s_{y}$) from Section 6.1, Table 4, into the exact expressions for $\theta$ (Eq. 11a), $\psi$ (Eq. 11b), $\varepsilon$ (Eq. 15a), and $\zeta$ (Eq. 16) given in Sections 3.3 and 3.4, we obtain expressions which do not have the form of a periodic series. Periodic series can be obtained instead by developing the angles’ expressions of Eqs. (14a, 14, 15a, 17b) around the dominant terms of the series ($\Delta n_{x},\Delta n_{y}$) and ($\Delta s_{x},\Delta s_{y}$), when they exist. If the first periodic term (with amplitudes $i_{1}$, $\theta_{1}$ and $\varepsilon_{1}$) dominates the other periodic terms ($i_{j}$, $\theta_{j}$ and $\varepsilon_{j}$ with $j>1$) and considering that the trends $tr_{x}$ and $tr_{y}$ are small, we can write the following developments for the inertial and orbital obliquities, the node longitude and the offset from the Cassini plane, up to order 2 in the small amplitudes of the other terms and in the trends:
@@
where we use the notation $\xi_{j}=f_{j}\,t+\varphi_{j}$. $i_{j}$, $\varepsilon_{j}$, $\theta_{j}$, $f_{j}$, and $\varphi_{j}$ are given in columns 3, 4, 6, and 7 in Table 4. As already noted after Eq. (15a), $\varepsilon$ is unaffected by the long-term trends. Developments of a higher order are usually necessary to reach a sufficient accuracy in the time domain.
@@
If the first, second, and possibly higher order terms are combined into pure trigonometric terms, each development can then be written as a trigonometric series:
@@
We do not show the trigonometric expansions here in their entirety, as they can include dozens and dozens of terms. For illustration purposes only, we provide truncated, and therefore imprecise, expansions in Tables 6 to 8, for the periodic and Poisson terms. At least as far as the most important terms are concerned, the $\theta$ and $\psi$ series, as well as the $\varepsilon$ and $\zeta$ series, tend to contain terms with the same arguments. The series for $\psi$ are multiplied by $\theta_{1}$ to obtain amplitudes of the same order as those of $\theta$. The Poisson series for the angles $\psi$, $\theta$ and $\zeta$ are plotted on Fig. 6. Between 2030 and 2035, during the JUICE mission timeframe, Ganymede’s Poisson terms can reach approximately up to $0.001^{\circ}$, $1^{\circ}$, and $0.0005^{\circ}$ for $\theta$, $\psi$, and $\zeta$, respectively. The periodic series of the orbital obliquity for Callisto closely matches the solution of Noyelles [2009] (see his Table 10), which was derived through frequency decomposition of a numerical integration of the Hamiltonian equations. Differences in amplitudes and frequencies are generally within a few percent, and are probably partly due to the use of a different ephemeris. This comparison of our analytical results and this numerical series provides support for our approach. A key advantage of our method is its ability to clearly identify the various contributing terms, including the Poisson terms.
@@
For Io and Europa, $\mathcal{T}_{cst}\sim\theta_{1}$ and $\mathcal{E}_{cst}\sim\varepsilon_{1}$, while most of their periodic terms are such that $|\mathcal{T}_{j}|\sim|\theta_{1}\mathcal{P}_{j}|\sim|\theta_{j}|$ as can be understood from the first-order term in Eq. (51a). Note, however, that the argument of each term is now a combination of two distinct arguments, rather than the original single argument of Table 4, so the periodicity will be different from the $s_{x}/s_{y}$ series. Most of their amplitudes $\mathcal{E}_{j}$ are also obtained essentially as first-order terms in Eq. (51c) and therefore tend to be similar to the amplitudes $\varepsilon_{j}$, but are not organized in the same order. For those first order terms, $\mathcal{Z}_{j}$ are not expected to be similar to $\varepsilon_{j}$, however, see Eq. (51d). For Ganymede (Callisto), because of the relatively large amplitudes of the second and third terms (second term) in Table 4, the amplitudes in Tables 7 and 6 do not necessarily correspond to those in Table 4 (especially for $\varepsilon$ in the case of Ganymede) and there are many high-order terms involved.
@@
With the exception of the orbital obliquity of Ganymede, the series of $\theta$, $\psi$, $\varepsilon$, and $\zeta$ help us to understand the temporal behavior of these angles (Figs. 2, 3, 4, and 5). The second terms of the series perturb the inertial obliquity of Io, Europa, Ganymede and Callisto by $26\%$, $5\%$, $22\%$, and $16\%$, respectively. The third term of Ganymede’s series perturbs its inertial obliquity by $18\%$, see third panel of Fig. 3. The main perturbation of $\varepsilon$ comes from $\mathcal{E}_{1}$ which is similar (or at least mainly originates from) to $\varepsilon_{4}$, $\varepsilon_{5}$, and $\varepsilon_{2}$ for Io, Europa, and Callisto, respectively. Europa’s orbital obliquity varies by $3\%$ over time, while its spin axis describes an almost circular trajectory in space, see Fig. 2. For Io and Callisto, $\varepsilon$ varies by $30\%$ and $55\%$, respectively. For Ganymede, $\varepsilon$ varies by almost $100\%$ around its mean value, between 0 and $0.6^{\circ}$, but this is poorly described by the $\varepsilon$ series. Callisto has the largest offset of the four satellites. Its periodic variations are dominated by the first term of the offset series with an amplitude $\mathcal{Z}_{1}=0.11^{\circ}$.
@@ 6.3 Series for $(\alpha_{S},\delta_{S})$ and $W$
The mean value of $\alpha_{S}$ is that of the Laplace plane ($\alpha_{LP}$), as already noticed in Eq. (7), see Section 3. For $\delta_{S}$, the mean value is the sum of $\delta_{LP}$ and of $\delta_{S\,0}\sim\theta_{1}^{2}\tan\delta_{LP}/4$, the latter being a small increment due to the power reduction of a squared sine term in the second-order part of Eq. (8b). As part of the transformation, this incremental quantity depends on the assumed interior model for the satellites (for instance, entirely solid or not). For $W$, we reexpress Eq. (31) assuming the following form:
@@
Using the transformations of Eqs. (8), (10), and (23) on the spin model developed in Section 6.1, we find values for the rates and for the amplitudes of the periodic series in $\alpha$, $\delta$ and $\mu$, as well as for the increment $\delta_{S\,0}$, see Tables 3 and 9. $\delta_{S\,0}$ ranges from $10^{-5}$ deg for Io to $2.5\,10^{-3}$ deg for Europa, see Table 3, which have respectively the smallest and largest dominant inertial obliquity amplitude $\theta_{1}$. Figure 7 shows the temporal evolution of the right ascension $\alpha_{S}$ and declination $\delta_{S}$ angles, whereas the evolution of $\mu$ is presented in Fig. 8.
@@
We now compare in the time domain the approximate series of Table 9 for $\alpha_{S}$ and $\delta_{S}$ to their exact transformation from Table 4 and Eqs. (5a)-(5b). Depending on whether or not the second-order transformation terms of Eqs. (8) are used or not, the error is generally less than $0.0002^{\circ}$ or $0.02^{\circ}$, respectively, see Table 10 for the Galilean satellites assumed to be entirely solid. For $\mu=\beta-\psi$, the comparison between the approximate series of Table 9 and the exact transformation of Eqs. (11b, 21) indicates an accuracy generally better than $0.0002^{\circ}$ or $0.013^{\circ}$ for the second or first-order transformations, respectively, see Table 10. Compared with a purely numerical solution obtained from ephemerides, these transformation errors should be added to the truncation errors in the original series in $n/s$, see Table 5, to obtain the total error budget. Neglecting secular terms would lead to typical errors of about $007^{\circ}, 0.0$$003^{\circ}$, and $0.006^{\circ}$ after $100$ years in $\alpha_{S}$, $\delta_{S}$, and $\mu$, respectively, larger than the combined truncation and second-order transformation errors.
Starting from the periodic series $\Delta\alpha_{S}$ and $\Delta\delta_{S}$ and Eqs. (14b, LABEL:eq_psiad, 15b, 17b), it is possible to obtain periodic series for the angles $\theta$, $\varepsilon$, $\psi$ and $\zeta$. We do not carry out this exercise here, as it would be very similar to the one based on the series $\Delta n_{x/y}$ and $\Delta s_{s/y}$, see Section 6.2.
@@
7 Recommendations for updating the IAU WGCCRE model
@@
Table 11: Rotation model of the IAU WGCCRE for the four Galilean satellites [Archinal et al., 2018]. The IAU WG arguments $J_{j}$ are expressed in the form $f_{j}\,T+\varphi_{j}$ in Eqs. (53-54), with $f_{j}$ and $\varphi_{j}$ the frequency and phase, respectively. $\alpha^{IAU}$ and $W^{IAU}$ are expressed as sine series whereas $\delta^{IAU}$ is expressed as a cosine series. The amplitudes are given in degrees and $T$ ($d$) is the time in Julian centuries (days) from J2000.
Since its very first report [Davies et al., 1980], the WGCCRE has openly assumed that the North pole of rotation of the Galilean satellites lies along the normal to their orbit, neglecting their small orbital obliquities in the expression for $\alpha^{IAU}$ and $\delta^{IAU}$. They also neglect the mean longitude variations and small librations in the periodic series of $W^{IAU}$. It can be indeed verified numerically that in the IAU WG solution, $\Delta W$ consists only in $\Delta\mu$, a projection of the $\Delta\alpha$ series, and assumes $\gamma_{u}=\Delta\mathcal{L}+\gamma_{f}=0$, see Eqs. (23e and 33). In addition to neglecting orbital obliquity and librations, the IAU WG model is also outdated. In particular, the periodic series based on Lieske ephemerides [Lieske, 1979] have remained unchanged ever since.
@@
Table 12: Proposed update for the solution ($\alpha_{S}$, $\delta_{S}$ and $W$) of the Galilean satellites considered as entirely rigid and using the JUP387 orbital theory, with a limited number of terms ($5$, 99, 77, and $4$, respectively) and a limited set of arguments ($14$). The arguments $\mathcal{J}_{j}$ are expressed in the form $f_{j}\,T+\varphi_{j}$ in Eqs. (53-54), with $f_{j}$ and $\varphi_{j}$ the frequency and phase, respectively. $\alpha_{S}$ and $W_{S}$ are expressed as sine series whereas $\delta_{S}$ is expressed as a cosine series. The amplitudes are given in degrees and $T$ ($d$) is the time in Julian centuries (days) from J2000. The column header shows the identification to the fundamental arguments, using the reduced set of arguments $\mathcal{J_{j}}$, whose physical origin is recalled on the first line. Compared to the full solution of Table 9 for the periodic terms, the precision is degraded to about $0.001^{\circ}$. Our model does not include any physical libration ($\gamma_{u}=0$).
@@
The Earth equatorial coordinates $\alpha_{LP}$ and $\delta_{LP}$ of the Laplace planes of our solution are close to those of the IAU WG solutions (they differ at most by $0.01^{\circ}$ and $0.005^{\circ}$, respectively), but more precise, see Table 1. We therefore recommend updating the mean values of $\alpha^{IAU}$ and $\delta^{IAU}$ with $\alpha_{LP}$ and $\delta_{LP}+\delta_{S\,0}$ as given in Tables 1 and 3, using at least three digits after the decimal point to take advantage of the improved precision achieved by recent ephemerides. The mean value for $\delta$ should also include the small increment $\delta_{S\,0}$, see Eq. (53b), because of the power reduction of a squared sine term in the second-order part of Eq. (8b). Computing a new independent $W_{0}$ requires control network computations which are beyond the scope of the present study. Here, the numerical value of $W_{0}$ is chosen to ensure that our defined prime meridian location ($W+\alpha_{S}\sin\delta_{S}$, where the $\alpha\sin\delta$ term accounts for the equator precession) matches that of the IAU WG model at the J2000 epoch. The corresponding corrections are 0.25-0.25, $2.8$, 4.6-4.6, and 11.5-11.5 km, respectively.
@@
The IAU WG rates in $\alpha^{IAU}$ ($9\,10^{-5}$ deg/y) and $\delta^{IAU}$ ($3\,10^{-5}$ deg/y) were last updated to match those of Jupiter in the 2005 report [Seidelmann et al., 2007]. Here, with the JUP387 ephemerides, we find rates in $\alpha_{S}$ and $\delta_{S}$ (e.g. $6.5\,10^{-5}$ deg/y and $2.4\,10^{-5}$ deg/y for Io, see Table 3) that are in good agreement with the rates in Jupiter’s equatorial coordinates of the current IAU WG solution ($6.499\,10^{-5}$ deg/y and $2.413\,10^{-5}$ deg/y). We therefore recommend updating the $\alpha$ and $\delta$ IAU WG rates for the Galilean satellites to the currently adopted rate for Jupiter. For the rates in $W$, those of Table 3, based on recent ephemerides, could replace the IAU WG rates.
@@
Each term in the periodic series of the IAU WG $(\alpha,\delta,W)$ solution can be replaced by the equivalent term of our solution for a solid satellite, see Table 9, which has (almost) the same period as the IAU WG term, but a different amplitude, which, as we will see later, ultimately leads to the neglect of one of the replaced terms (the term for Callisto at period 5.93 years). Thanks to the use of more recent and precise ephemeris, the second-order part of the transformation to $\alpha$ and $\delta$, and the non-zero obliquity, we find several additional terms that can be considered for an update of the IAU WG model, depending on the precision required. For Callisto, three of these additional terms have amplitudes larger than the sixth term of the series, which corresponds to the third term of the IAU WG series (the very one we will be neglecting below). Note that when considering different interior structures for a satellite with a global ocean (see for example section 8.2), some amplitudes may be significantly different from their solid counterparts if the forcing frequency is close enough to one of the free mode frequencies, so that the order of the different terms may differ.
@@
The IAU WG arguments are characterized by positive rates while in our solution (Table 9), the rates of the arguments are mostly negative, which makes more sense from a dynamic point of view. Throughout our successive developments, from ephemerides to $\Delta\alpha_{S},\Delta\delta_{S}$ and $\Delta\mu$ series, we have kept, at least for the first-order terms of the transformation, signs of rates that correspond to the convention used for the periodic series of Eq. (48) for $n_{x/y}$, where a retrograde precession of the orbital node is characterized by a negative rate. This convention allows for the use of the dynamical Eq. (13) with $\dot{\Omega}=f_{j}$ to get the amplitudes (Eq. 49) of the periodic series of Eq. (50) for $s_{x/y}$. In this way, we can easily compare the forced frequencies $f_{j}$ with the $\kappa/C$ free frequency to identify the potential for resonant amplification. With this convention, the phases remain unchanged by the transformation from $\Delta s_{x}/\Delta s_{y}$ to $\Delta\alpha_{S}/\Delta\delta_{S}/\Delta\mu$ and, as a result, differ from those of the IAU WG arguments, while the different terms of $\Delta\alpha_{S}$ and $\Delta\delta_{S}$ have amplitude of opposite signs for the same argument, when the signs are the same in the IAU WG series.

### v2

@@ 6.2 Series for obliquity, node longitude, and offset
The temporal evolution of the inertial ($\theta$, Eq. 12a) and orbital ($\varepsilon$, Eq. 14) obliquities, node longitude ($\psi$, Eq. 12b), and offset ($\zeta$, Eq. 17) obtained from the series of Table 3 and Table 4 are shown in Figs. 4, 5, and 6. We see that the deviation from the Cassini plane is small for Europa (usually less than $0.006^{\circ}$). For comparison with the spin Euler angles, the inclination ($i$, Eq. 13a) and orbital node longitude ($\Omega$, Eq. 13b) are also shown in Fig. 4 and 5, respectively. The typical precision (or truncation error) of the time-series $\varepsilon(t)$ and $\theta(t)$ over the 1900-2100 interval, reconstructed from Table 4, is by construction similar to that of the time-series $s_{x/y}(t)$, that is approximately $0.0003^{\circ}$ (see Table 5). This corresponds to a relative error below $1\%$, except in the case of Io’s orbital obliquity which is particularly small. In contrast, the longitude of the ascending node is a circulating angle, and the error associated with its reconstruction using Table 4 and Eq. (12b) is significantly larger but remains below $0.6^{\circ}$ and therefore below the percent level. The absolute precision on the reconstructed $\zeta(t)$ is similar to that of $\varepsilon(t)$, but its relative precision may be larger than that of $\varepsilon(t)$ when the offset is small (e.g. Europa, $5\%$). As for $s_{x/y}(t)$, neglecting terms related to trends would lead after $100$ years to typical errors larger than the truncation error in $\theta$, $\psi$, and $\zeta$ ($\varepsilon$ is not affected by the trends).
When we inject the series ($\Delta n_{x},\Delta n_{y}$) and ($\Delta s_{x},\Delta s_{y}$) from Section 6.1, Table 4, into the exact expressions for $\theta$ (Eq. 12a), $\psi$ (Eq. 12b), $\varepsilon$ (Eq. 16b), and $\zeta$ (Eq. 17) given in Sections 3.3 and 3.4, we obtain expressions which do not have the form of a periodic series. Periodic series can be obtained instead by developing the angles’ expressions of Eqs. (15a, 15, 16b, 18b) around the dominant terms of the series ($\Delta n_{x},\Delta n_{y}$) and ($\Delta s_{x},\Delta s_{y}$), when they exist. If the first periodic term (with amplitudes $i_{1}$, $\theta_{1}$ and $\varepsilon_{1}$) dominates the other periodic terms ($i_{j}$, $\theta_{j}$ and $\varepsilon_{j}$ with $j>1$) and considering that the trends $tr_{x}$ and $tr_{y}$ are small, we can write the following developments for the inertial and orbital obliquities, the node longitude and the offset from the Cassini plane, up to order 2 in the small amplitudes of the other terms and in the trends:
@@
where we use the notation $\xi_{j}=f_{j}\,t+\varphi_{j}$. $i_{j}$, $\varepsilon_{j}$, $\theta_{j}$, $f_{j}$, and $\varphi_{j}$ are given in columns 2, 3, 4, 6, and 7 in Table 4. As already noted after Eq. (16b), $\varepsilon$ is unaffected by the long-term trends. Developments of a higher order are usually necessary to reach a sufficient accuracy in the time domain.
@@
Rewriting products of trigonometric functions in terms of trigonometric functions with combined arguments, each development can then be written as a trigonometric series:
@@
We do not show the trigonometric expansions here in their entirety, as they can include dozens and dozens of terms. For illustration purposes only, we provide truncated, and therefore imprecise, expansions in Tables 6 to 8, for the periodic and Poisson terms. At least as far as the most important terms are concerned, the $\theta$ and $\psi$ series, as well as the $\varepsilon$ and $\zeta$ series, tend to contain terms with the same arguments. The series for $\psi$ are multiplied by $\theta_{1}$ to obtain amplitudes of the same order as those of $\theta$. The Poisson series for the angles $\psi$, $\theta$ and $\zeta$ are plotted on Fig. 7. Between 2030 and 2035, during the Juice mission timeframe, Ganymede’s Poisson terms can reach approximately up to $0.001^{\circ}$, $1^{\circ}$, and $0.0005^{\circ}$ for $\theta$, $\psi$, and $\zeta$, respectively. The periodic series of the orbital obliquity for Callisto closely matches the solution of Noyelles [2009] (see his Table 10), which was derived through frequency decomposition of a numerical integration of the Hamiltonian equations. Differences in amplitudes and frequencies are generally within a few percent, and are probably partly due to the use of a different ephemeris. This comparison of our analytical results and this numerical series provides support for our approach. A key advantage of our method is its ability to clearly identify the various contributing terms, including the Poisson terms in $\theta$, $\psi$ and $\zeta$, see Eq. (53).
@@
For Io and Europa, $\mathcal{T}_{cst}\sim\theta_{1}$ and $\mathcal{E}_{cst}\sim\varepsilon_{1}$, while most of their periodic terms are such that $|\mathcal{T}_{j}|\sim|\theta_{1}\mathcal{P}_{j}|\sim|\theta_{j}|$ as can be understood from the first-order term in Eq. (52a). Note, however, that the argument of each term is now a combination of two distinct arguments, rather than the original single argument of Table 4, so the periodicity will be different from the $s_{x}/s_{y}$ series. Most of their amplitudes $\mathcal{E}_{j}$ are also obtained essentially as first-order terms in Eq. (52c) and therefore tend to be similar to the amplitudes $\varepsilon_{j}$, but are not organized in the same order. For those first order terms, $\mathcal{Z}_{j}$ are not necessarily expected to have an amplitude similar to that of $\varepsilon_{j}$, however, see Eq. (52d). For Ganymede (Callisto), because of the relatively large amplitudes of the second and third terms (second term) in Table 4, the amplitudes in Tables 7 and 6 do not necessarily correspond to those in Table 4 (especially for $\varepsilon$ in the case of Ganymede) and there are many high-order terms involved.
@@
With the exception of the orbital obliquity of Ganymede, the series of $\theta$, $\psi$, $\varepsilon$, and $\zeta$ help us to understand the temporal behavior of these angles (Figs. 3, 4, 5, and 6). The second terms of the series perturb the inertial obliquity of Io, Europa, Ganymede and Callisto by $26\%$, $5\%$, $22\%$, and $16\%$, respectively. The third term of Ganymede’s series perturbs its inertial obliquity by $18\%$, see third panel of Fig. 4. The main perturbation of $\varepsilon$ comes from $\mathcal{E}_{1}$ which is similar (or at least mainly originates from) to $\varepsilon_{4}$, $\varepsilon_{5}$, and $\varepsilon_{2}$ for Io, Europa, and Callisto, respectively. Europa’s orbital obliquity varies by $3\%$ over time, while its spin axis describes an almost circular trajectory in space, see Fig. 3. For Io and Callisto, $\varepsilon$ varies by $30\%$ and $55\%$, respectively. For Ganymede, $\varepsilon$ varies by almost $100\%$ around its mean value, between 0 and $0.6^{\circ}$, but this is poorly described by the $\varepsilon$ series. Callisto has the largest offset of the four satellites. Its periodic variations are dominated by the first term of the offset series with an amplitude $\mathcal{Z}_{1}=0.11^{\circ}$.
@@ 6.3 Series for $(\alpha_{S},\delta_{S})$ and $W$
The mean value of $\alpha_{S}$ is that of the Laplace plane ($\alpha_{LP}$), as already noticed in Eq. (8), see Section 3. For $\delta_{S}$, the mean value is the sum of $\delta_{LP}$ and of $\delta_{S\,0}\sim\theta_{1}^{2}\tan\delta_{LP}/4$, the latter being a small increment due to the constant term induced by the power reduction of a squared sine term in the second-order part of Eq. (9b). As part of the transformation, this incremental quantity depends on the assumed interior model for the satellites (for instance, entirely solid or not). For $W$, we reexpress Eq. (32) assuming the following form:
@@
Using the transformations of Eqs. (9), (11), and (24) on the spin model developed in Section 6.1, we find values for the rates and for the amplitudes of the periodic series in $\alpha$, $\delta$ and $\mu$, as well as for the increment $\delta_{S\,0}$, see Tables 3 and 9. $\delta_{S\,0}$ ranges from $10^{-5}$ deg for Io to $2.5\times 10^{-3}$ deg for Europa, see Table 3, which have respectively the smallest and largest dominant inertial obliquity amplitude $\theta_{1}$. Figure 8 shows the temporal evolution of the right ascension $\alpha_{S}$ and declination $\delta_{S}$ angles, whereas the evolution of $\mu$ is presented in Fig. 9.
@@
We now compare in the time domain the approximate series of Table 9 for $\alpha_{S}$ and $\delta_{S}$ to their exact transformation from Table 4 and Eqs. (6a)-(6b). Depending on whether or not the second-order transformation terms of Eqs. (9) are used or not, the error is generally less than $0.0002^{\circ}$ or $0.02^{\circ}$, respectively, see Table 10 for the Galilean satellites assumed to be entirely solid. For $\mu=\beta-\psi$, the comparison between the approximate series of Table 9 and the exact transformation of Eqs. (12b, 22) indicates an accuracy generally better than $0.0002^{\circ}$ or $0.013^{\circ}$ for the second or first-order transformations, respectively, see Table 10. Compared with a purely numerical solution obtained from ephemerides, these transformation errors should be added to the truncation errors in the original series in $n/s$, see Table 5, to obtain the total error budget. Neglecting secular terms would lead to typical errors of about $007^{\circ}, 0.0$$003^{\circ}$, and $0.006^{\circ}$ after $100$ years in $\alpha_{S}$, $\delta_{S}$, and $\mu$, respectively, larger than the combined truncation and second-order transformation errors.
Starting from the periodic series $\Delta\alpha_{S}$ and $\Delta\delta_{S}$ and Eqs. (15b, LABEL:eq_psiad, 16b, 18b), it is possible to obtain periodic series for the angles $\theta$, $\varepsilon$, $\psi$ and $\zeta$. We do not carry out this exercise here, as it would be very similar to the one based on the series $\Delta n_{x/y}$ and $\Delta s_{s/y}$, see Section 6.2.
@@
7 Recommendations for updating the IAU WGCCRE solution
@@
Table 11: Rotation solution of the IAU WGCCRE for the four Galilean satellites [Archinal et al., 2018]. The IAU WG arguments $J_{j}$ are expressed in the form $f_{j}\,T+\varphi_{j}$ in Eqs. (54-55), with $f_{j}$ and $\varphi_{j}$ the frequency and phase, respectively. $\alpha^{IAU}$ and $W^{IAU}$ are expressed as sine series whereas $\delta^{IAU}$ is expressed as a cosine series. The amplitudes are given in degrees and $T$ ($d$) is the time in Julian centuries (days) from J2000.
Since its very first report [Davies et al., 1980], the WGCCRE has openly assumed that the North pole of rotation of the Galilean satellites lies along the normal to their orbit, neglecting their small orbital obliquities in the expression for $\alpha^{IAU}$ and $\delta^{IAU}$. They also neglect the mean longitude variations and small librations in the periodic series of $W^{IAU}$. It can be indeed verified numerically that in the IAU WG solution, $\Delta W$ consists only in $\Delta\mu$, a projection of the $\Delta\alpha$ series, and assumes $\gamma_{u}=\Delta\mathcal{L}+\gamma_{f}=0$, see Eqs. (24e and 34). In addition to neglecting orbital obliquity and librations, the IAU WG model is also outdated. In particular, the periodic series based on Lieske ephemerides [Lieske, 1979] have remained unchanged ever since.
@@
Table 12: Proposed update for the solution ($\alpha_{S}$, $\delta_{S}$ and $W$) of the Galilean satellites considered as entirely rigid and using the JUP387 orbital theory, with a limited number of terms ($5$, 99, 77, and $4$, respectively) and a limited set of arguments ($14$). The arguments $\mathcal{J}_{j}$ are expressed in the form $f_{j}\,T+\varphi_{j}$ in Eqs. (54-55), with $f_{j}$ and $\varphi_{j}$ the frequency and phase, respectively. $\alpha_{S}$ and $W_{S}$ are expressed as sine series whereas $\delta_{S}$ is expressed as a cosine series. The amplitudes are given in degrees and $T$ ($d$) is the time in Julian centuries (days) from J2000. The column header shows the identification to the fundamental arguments, using the reduced set of arguments $\mathcal{J_{j}}$, whose physical origin is recalled on the first line. Compared to the full solution of Table 9 for the periodic terms, the precision is degraded to about $0.001^{\circ}$. Our model does not include any physical libration ($\gamma_{u}=0$).
@@
The Earth equatorial coordinates $\alpha_{LP}$ and $\delta_{LP}$ of the Laplace planes of our solution are close to those of the IAU WG solutions (they differ at most by $0.01^{\circ}$ and $0.005^{\circ}$, respectively), but more precise, see Table 1. We therefore recommend updating the mean values of $\alpha^{IAU}$ and $\delta^{IAU}$ with $\alpha_{LP}$ and $\delta_{LP}+\delta_{S\,0}$ as given in Tables 1 and 3, using at least three digits after the decimal point to take advantage of the improved precision achieved by recent ephemerides. The mean value for $\delta$ should also include the small increment $\delta_{S\,0}$, see Eq. (54b), because of the power reduction of a squared sine term in the second-order part of Eq. (9b). Computing a new independent $W_{0}$ requires control network computations which are beyond the scope of the present study. Here, the numerical value of $W_{0}$ is chosen to ensure that our defined prime meridian location ($W+\alpha_{S}\sin\delta_{S}$, where the $\alpha_{S}\sin\delta_{S}$ term accounts for the equator precession) matches that of the IAU WG model at the J2000 epoch. The corresponding corrections are 0.25-0.25, $2.8$, 4.6-4.6, and 11.5-11.5 km, respectively.
@@
Originally, the IAU WG adopted the Jovian $\alpha$ and $\delta$ rates ($9\times 10^{-5}$ deg/y and $3\times 10^{-5}$ deg/y in $\alpha$ and $\delta$, respectively) from Lieske [1979] and used them as the rates for the Galilean satellites as well. It was not until the 2005 report (see Seidelmann et al. 2007) that the Jovian rates were updated, but the IAU WG did not update those of the satellites at the same time. Here, with the JUP387 ephemerides, we find rates in $\alpha_{S}$ and $\delta_{S}$ (e.g. $6.5\times 10^{-5}$ deg/y and $2.4\times 10^{-5}$ deg/y for Io, see Table 3) that are in good agreement with the rates in Jupiter’s equatorial coordinates of the current IAU WG solution ($6.499\times 10^{-5}$ deg/y and $2.413\times 10^{-5}$ deg/y). We therefore recommend updating the $\alpha$ and $\delta$ IAU WG rates for the Galilean satellites to the currently adopted rate for Jupiter. For the rates in $W$, those of Table 3, based on recent ephemerides, could replace the IAU WG rates.
@@
Each term in the periodic series of the IAU WG $(\alpha,\delta,W)$ solution can be replaced by the equivalent term of our solution for a solid satellite, see Table 9, which has (almost) the same period as the IAU WG term, but a different amplitude, which, as we will see later, ultimately leads to the neglect of one of the replaced terms (the term for Callisto at period 5.93 years). Thanks to the use of more recent and precise ephemeris, the second-order part of the transformation to $\alpha$ and $\delta$, and the non-zero obliquity, we find several additional terms that can be considered for an update of the IAU WG solution, depending on the precision required. For Callisto, three of these additional terms have amplitudes larger than the sixth term of the series, which corresponds to the third term of the IAU WG series (the very one we will be neglecting below). Note that when considering different interior structures for a satellite with a global ocean (see for example section 8.2), some amplitudes may be significantly different from their solid counterparts if the forcing frequency is close enough to one of the free mode frequencies, so that the order of the different terms may differ.
@@
The IAU WG arguments are characterized by positive rates while in our solution (Table 9), the rates of the arguments are mostly negative, which makes more sense from a dynamic point of view (precessions are usually retrograde motions). Throughout our successive developments, from ephemerides to $\Delta\alpha_{S},\Delta\delta_{S}$ and $\Delta\mu$ series, we have kept, at least for the first-order terms of the transformation, signs of rates that correspond to the convention used for the periodic series of Eq. (49) for $n_{x/y}$, where a retrograde precession of the orbital node is characterized by a negative rate. This convention allows for the use of the dynamical Eq. (14) with $\dot{\Omega}=f_{j}$ to get the amplitudes (Eq. 50) of the periodic series of Eq. (51) for $s_{x/y}$. In this way, we can easily compare the forced frequencies $f_{j}$ with the $\kappa/C$ free frequency to identify the potential for resonant amplification. With this convention, the phases remain unchanged by the transformation from $\Delta s_{x}/\Delta s_{y}$ to $\Delta\alpha_{S}/\Delta\delta_{S}/\Delta\mu$ and, as a result, differ from those of the IAU WG arguments, while the different terms of $\Delta\alpha_{S}$ and $\Delta\delta_{S}$ have amplitude of opposite signs for the same argument, when the signs are the same in the IAU WG series.